OpenStudy (precal):

Find two values on the interval (0, 2pi) where the slope of the tangent to the graph of f(x)=cos(2x) is equal to square root of 3

OpenStudy (anonymous):

take the derivative set it equal to \(\sqrt3\) solve for \(x\)

OpenStudy (precal):

I found the derivative f ' (x) =-2sin(2x)

OpenStudy (precal):

I am stuck on solving for x

OpenStudy (anonymous):

k that looks good

OpenStudy (anonymous):

divide by \(-2\) then think or appeal to the unit circle

OpenStudy (precal):

\[-2\sin(2x)=\sqrt{3}\] \[\sin(2x)=\frac{ \sqrt{3} }{ -2 }\]

OpenStudy (anonymous):

right

OpenStudy (precal):

what about the 2x?

OpenStudy (anonymous):

should be familiar

OpenStudy (anonymous):

you know a number where the sine is \(-\frac{\sqrt3}{2}\) ?

OpenStudy (precal):

at 120 degrees or (2pi)3

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

you want to be in quadrant 3 or 4 because sine is negative

OpenStudy (precal):

sorry, let me go look at a unit circle

OpenStudy (anonymous):

k

OpenStudy (anonymous):

use radians, not degrees none of this works in degrees

OpenStudy (anonymous):

let me know when you get \(\frac{4\pi}{3}\)

OpenStudy (anonymous):

also \(\frac{5\pi}{3}\)

OpenStudy (precal):

\[x =\frac{ 4\pi }{ 3 } \] \[x=\frac{ 5\pi }{ 3}\]

OpenStudy (anonymous):

ok good but you have \(2x\) not \(x\)

OpenStudy (precal):

sorry I had to look it up and type them yes sorry and then I solve for x

OpenStudy (anonymous):

so it should be \(2x=\frac{4\pi}{3}\iff x=\frac{4\pi}{6}=\frac{2\pi}{3}\) in other words, divide by two

OpenStudy (precal):

any hints so that I can recall this information without looking it up everytime, I only have the first quadrant memorize

OpenStudy (precal):

and x = (5pi)/6

OpenStudy (anonymous):

similarly \[2x=\frac{5\pi}{3}\iff x=\frac{5\pi}{6}\] the solving for \(x\) part with \(2x\) is not supposed to be the hard part

OpenStudy (precal):

true, sorry, I am a little slow

OpenStudy (anonymous):

no hints, just visualize it if sine is negative you are below the x axis, not above

OpenStudy (anonymous):

a little slow? i doubt it btw this is the first time i looked "anne fibian" great!

OpenStudy (precal):

Yes, old joke I found in a textbook

OpenStudy (precal):

there is a great little exercise for collecting data and creating the sine and cosine graphs using a frog

OpenStudy (anonymous):

?

OpenStudy (precal):

just helps me remember where the sine graph begins

OpenStudy (anonymous):

got it

OpenStudy (precal):

just explaining why I use the username Precal and the frog "Anne Fibian"

OpenStudy (precal):

Thanks so much, you give me hope that one day I too can conquer Calculus